OpenStudy (anonymous):

Compute the probability of X successes, using table B in Appendix C in the text book. n = 15, p = 0.8, X=10 Choose one answer. a. 0.188 b. 0.103 c. 0.206 d. 0.010

OpenStudy (anonymous):

the Binomial Probability distribution = \[\left(\begin{matrix}n \\ x\end{matrix}\right)\] *(p)^x*(1-p)^(n-x) where n = number of samples (15) x = number of successes (10) p = probability of success (.8) therefore: \[\left(\begin{matrix}15 \\ 10\end{matrix}\right)\] *(.8)^10*(1-.8)^(15-10) = .103